How many liters of 10% alcohol solution and 5% alcohol solution must be mixed to get 40 liters of 8%?

Found 2 solutions by stanbon, ikleyn:
Answer by stanbon(75887)

How many liters of 10% alcohol solution and 5% alcohol solution must be mixed to get 40 liters of 8%?
----
Equation:
alcohol + alcohol = alcohol
-------
0.10x + 0.05(40-x) = 0.08*40
--------
10x + 5*40 - 5x = 8*40
--------
5x = 3*40
x = 24 liters (amt. of 10% solution needed)
---
40-x = 16 liters (amt. of 5% solution needed)
--------------
Cheers,
Stan H.
-----------
----

Answer by ikleyn(52781)

.
Similar problem was solved in the lesson  Solving typical word problems on mixtures for solutions  in this site  (Problem 1).

Read it attentively and then solve your problem by substituting your data.